# Surface groups among cubulated hyperbolic and one-relator groups

### v1

@@ Abstract
Let $X$ be a non-positively curved cube complex with hyperbolic fundamental group. We prove that $\pi_{1}(X)$ has a non-free subgroup of infinite index unless $\pi_{1}(X)$ is either free or a surface group, answering a question of Wise. A similar result for one-relator groups follows, answering a question posed by several authors. The proof relies on a careful analysis of free and cyclic splittings of cubulated groups.
Surface groups – the fundamental groups of closed, aspherical surfaces – and free groups are often seen as close cousins. One way this can be made precise is to look at their subgroup structures. The Nielsen–Schreier theorem asserts that every subgroup of a free group is free, while Whitehead also proved that every subgroup of infinite index in a surface group is free. It is natural to wonder whether this property characterises free and surface groups. Conjectures and questions along these lines have been stated under various hypotheses [4, 3, 20, 18, 55].
@@
In particular, Theorem A applies to $C^{\prime}(1/6)$ small-cancellation groups [54].
@@
If $G$ is hyperbolic then the result follows from Theorem A. Otherwise, Caprace–Haglund proved that $G$ contains a cubically convex-cocompact subgroup isomorphic to $\mathbb{Z}^{2}$ [10, Corollary 4], and the result follows unless $\mathbb{Z}^{2}$ has finite index. But, in the latter case, $G$ is a surface group by Bieberbach’s theorem. ∎
This corollary answers a question asked by Wise in the 2014 ICM proceedings [55, Question 13.50].
@@
Historically, Question 0.1 has received the most attention when $G$ is a one-relator group. In this context, the question was attributed to Fine–Rosenberger–Wienke in the Kourovka notebook [18, Question 20.22], but it has also received attention in many other works on one-relator groups, sometimes with additional hypotheses [3, 4, 14, 19]. In this context, it is sometimes referred to as the surface group conjecture. In [52], the author proved the surface group conjecture for Sela’s limit groups, and also for graphs of free groups with cyclic edge groups. [^2] The latter class includes the classes of cyclically pinched and conjugacy pinched one-relator groups considered in [14, 19].
Most recently, the surface group conjecture was addressed by Gardam–Kielak–Logan [20, Conjecture 1.4], who proved it in the case when $G$ has two generators. Combining their work with Theorem A and other recent advances in our understanding of one-relator groups, the surface group conjecture follows.
@@
In fact, the proof of Theorem D produces a finitely generated non-free subgroup unless $G$ has two generators. In the two-generator case, the stronger result is not true: every finitely generated subgroup of infinite index in the $(1,2)$-Baumslag–Solitar group $BS(1,2)$ is either trivial or isomorphic to $\mathbb{Z}$, but $BS(1,2)$ contains a subgroup isomorphic to the infinitely generated non-free group $\mathbb{Q}_{2}$ of dyadic rationals. (Cf. Question 5.4 below.)
@@
Assuming $\pi_{1}(X)$ is one-ended, our next task is to analyse splittings over $\mathbb{Z}$, and to do this we restrict to the case when $\pi_{1}(X)$ is hyperbolic. The approach generalises work of Cashen–Macura [13] and Cashen [12], who showed that a line pattern in a free group does not split over $\mathbb{Z}$ if and only if there is a bound on the diameters of cut pairs in the Whitehead graph. We prove an analogous result, Theorem 3.11, which asserts that $\pi_{1}(X)$ is cyclically indecomposable if and only if there is a uniform bound on the widths of 2-cuts. The point is that cyclic splittings correspond to periodic 2-cuts, from which a pigeonhole argument implies a bound on the width of a 2-cut if $\pi_{1}(X)$ is cyclically indecomposable. Using the projection-complex machinery of Bestvina–Bromberg–Fujiwara [5], we strengthen this to a uniform bound on the widths of $k$-cuts, independent of $k$.
@@
The paper is structured as follows. Section 1 contains preliminaries on non-positively curved complexes, cube complexes and simplicial complexes. Experts may wish to jump straight to Section 2, where we define Whitehead complexes and use them to analyse free splittings, including proving Theorem E. In Section 3 we extend the analysis to cyclic splittings in the hyperbolic case. Section 4 contains the proofs of the main theorems, while some open questions are listed in Section 5.
@@ Acknowledgements
I am grateful to Mark Hagen for useful conversations about this work, most especially for suggesting Lemma 3.16, which greatly simplifies the argument. Thanks also to Macarena Arenas, Jack Button, Chris Cashen, Mark Hagen again, Marco Linton, Lars Louder and Sam Shepherd for comments on an early draft.
@@ 1.1 Non-positively curved cube complxes
A cube complex $X$ is a cell complex constructed from a disjoint union of finite-dimensional unit cubes by identifying faces homeomorphically [8, Example I.7.40(4)].
The link $\mathrm{Lk}_{X}(x)$ of a point $x$ in a metric polyhedral complex $X$ is the space of directions at $x$, i.e. the space of germs of geodesics emanating from $x$ [8, I.7.38], and admits a natural angular metric. When $X$ is a cube complex and $x$ is a vertex, $\mathrm{Lk}_{X}(x)$ has the natural structure of a simplicial complex, where the simplices are the corners of the cubes incident at $x$.
As long as a cube complex $X$ is finite-dimensional, $X$ admits a natural complete geodesic metric in which the interior of every $k$-cube is locally isometric to $\mathbb{R}^{k}$. We will refer to this metric as the $\ell^{2}$-metric on $X$. Gromov’s link condition asserts that a polyhedral complex is non-positively curved – i.e. the metric is locally CAT(0) – if and only if the angular metric on the link of any point is CAT(1) [22, 4.2.A]. Gromov further proved that, when $X$ is a cube complex, the link condition is satisfied if and only if the link of every vertex is flag, meaning that, for every $k\geq 2$, every $k$-clique in the 1-skeleton of $X$ is the boundary of a $k$-simplex [22, 4.2.C].
The features of a non-positively curved cube complex $X$ can often be captured by studying their hyperplanes. Let $C$ be a $k$-dimensional cube, identified with $[-1/2,1/2]^{k}\subseteq\mathbb{R}^{k}$. A hypercube of $C$ is an intersection of $C$ with any coordinate hyperplane of $\mathbb{R}^{k}$. The set of hypercubes of cubes in $X$ form a natural directed system under inclusion, and gluing along these inclusions gives a cube complex with a natural map to $X$. The path components of this cube complexes are the hyperplanes of $X$.
@@
Finally, we can make use of the following important lemma [8, Proposition II.4.14].
@@ 1.2 CAT(0) cube complexes
When $X$ is non-positively curved, the natural geodesic metric on the universal cover $\widetilde{X}$ is CAT(0), by the Cartan–Hadamard theorem [22, §4]. This gives us access to the standard tools of CAT(0) geometry detailed in [8], which are especially useful when studying a closed, convex subspace $Y\subseteq\widetilde{X}$. For instance, we may define the orthogonal projection $\pi_{Y}:\widetilde{X}\to Y$ as in [8, Proposition II.2.4].
One important consequence of the uniqueness of geodesics in CAT(0) spaces is that any local isometry between CAT(0) spaces is a global isometric embedding [8, Proposition II.4.14]. Thus, the map from a hyperplane $H\to X$ lifts to a convex embedding of the universal cover $\widetilde{H}\hookrightarrow\widetilde{X}$, and so the hyperplanes of $\widetilde{X}$ are the universal covers of the hyperplanes of $X$.
@@
Fortunately, we may always assume that $X$ is essential by invoking a result of Caprace–Sageev [11, Proposition 3.5].
@@
Note that the $\ell^{1}$ metric on $X^{(1)}$ is also geodesic, and that the $\ell^{1}$ distance between two points is exactly the number of hyperplanes that separate them. The next lemma is useful for relating the $\ell^{1}$ and $\ell^{2}$ distances.
@@
Let $Y_{1},Y_{2}$ be convex subcomplexes of a CAT(0) cube complex $\widetilde{X}$. Suppose that a geodesic $\gamma$ minimises the distance from $Y_{1}$ to $Y_{2}$. A hyperplane $H$ separates $Y_{1}$ from $Y_{2}$ if and only if $\gamma$ crosses $H$.
@@
The necessity is obvious, so it remains to prove that, if $\gamma$ crosses $H$, then $H$ separates $Y_{1}$ from $Y_{2}$. Suppose that $\gamma(0)\in Y_{1}$ and $\gamma$ crosses $H$ at $\gamma(t)$. If $H$ intersects $Y_{1}$ then $\gamma(0)$, $\gamma(t)$ and $\pi_{H}(\gamma(0))$ forms a triangle with two right angles, contradicting the hypothesis that $\widetilde{X}$ is CAT(0). Therefore, $H$ is disjoint from $Y_{1}$, and likewise from $Y_{2}$. Since $\gamma$ crosses $H$ once, it follows that $H$ separates $Y_{1}$ from $Y_{2}$, as required. ∎
We can now state Haglund’s theorem [25, Theorem H]
@@
One useful consequence is to characterise geometrically well-behaved subgroups. A subgroup $H$ of the fundamental group of a non-positively curved cube complex $X$ is called cubically convex-cocompact if $H$ acts cocompactly on a subcomplex $Y\subseteq\widetilde{X}$ that is convex in the $\ell^{2}$ metric.
@@
Let $X$ be a compact, non-positively curved cube complex. A subgroup $H$ of $\pi_{1}(X)$ is cubically convex-cocompact if and only if the $H$-orbit of some vertex is quasiconvex in the $\ell^{1}$ metric.
@@ 1.3 Simplicial complexes
Let $A$ be a simplicial complex and $B\subseteq A$ a full subcomplex. The open star of $B$, $\mathrm{St}_{A}(B)$, is the union of the set of simplices of $A$ that contain a simplex of $B$ as a face. The link of $B$, $\mathrm{Lk}_{A}(B)$, is the set of faces of simplices of $\mathrm{St}_{A}(B)$ that are disjoint from $B$.
@@
In the specific case where $B$ is a set of $k$ pairwise non-adjacent vertices, we say that $B$ is a cut set for $A$. If #$\#B=1$ we say $B$ is a cut vertex, and if #$\#B=2$ we say $B$ is a cut pair. If $A$ has no cut set of cardinality at most $k$, then $A$ is said to be $k$-connected.
@@ 2 Free splittings
In 1936, J.H.C. Whitehead explained how to determine whether an element $w$ of a free group $F$ is primitive. For a fixed basis $B$ of $F$, his algorithm uses a graph $\mathrm{Wh}_{B}(w)$ which is now known as the Whitehead graph [51]. His algorithm can equally be interpreted as searching for free splittings $F\cong\langle w\rangle*F^{\prime}$. Many elaborations and generalisations of Whitehead’s argument have been given over the years by, among others, Lyon [33], Stallings [46] and Stong [47]. Diao–Feighn, building on work of Gersten [21], gave an algorithm to compute the Grushko decomposition of a graph of free groups [15]. [^3] See the recent preprint of Dicks for a comprehensive history of this topic [16, §2].
The power of this point of view is that global information about a group $G$ – whether or not $G$ is free, or splits freely, or most generally the structure of the Grushko decomposition of $G$ – is encoded in the local information of a Whitehead graph, which can be read off a presentation. The following theorem, which is the result of combining Whitehead’s algorithm with a theorem of Shenitzer [41, Theorem 5], exemplifies this idea in the case of one-relator groups.
@@ 2.1 Bounding hyperplanes and Whitehead complexes
The Helly property for hyperplanes of CAT(0) cube complexes, proved by Sageev, asserts that if a set of hyperplanes intersect pairwise then their intersection is non-empty [38, Theorem 4.14]. Hence, Whitehead complexes are flag.
@@
Whitehead complexes are useful because they provide a topological model for the complement of $Y$: we shall see that $Wh_{X}(Y)$ is homotopy equivalent to the complement $\widetilde{X}\smallsetminus Y$. In order to describe this homotopy equivalence, it is useful to define the link of $Y$, and to extend our discussion of bounding hyperplanes to codimension-$k$ hyperplanes.
@@
By definition, $\mathrm{Wh}_{\widetilde{X}}(Y)$ is the complex obtained by contracting each $K_{Y}$ factor to a point in the definition of $\mathrm{Lk}_{\widetilde{X}}(Y)$. Since $\mathrm{Lk}_{\widetilde{X}}(Y)$ is a cell complex and each $K_{Y}$ is contractible, the result follows by, for example, [27, Proposition 0.16]. ∎

### v2

@@ Abstract
Let $X$ be a non-positively curved cube complex with hyperbolic fundamental group. We prove that $\pi_{1}(X)$ has a non-free subgroup of infinite index unless $\pi_{1}(X)$ is either free or a surface group, answering questions of Gromov and Whyte (in a special case) and Wise. A similar result for one-relator groups follows, answering a question posed by several authors. The proof relies on a careful analysis of free and cyclic splittings of cubulated groups.
Surface groups – the fundamental groups of closed, aspherical surfaces – and free groups are often seen as close cousins. One way this can be made precise is to look at their subgroup structures. The Nielsen–Schreier theorem asserts that every subgroup of a free group is free, while Johansson also proved that every subgroup of infinite index in a surface group is free [34]. It is natural to wonder whether this property characterises free and surface groups. Conjectures and questions along these lines have been stated under various hypotheses [3, 4, 5, 23, 35, 68].
@@
This answers questions of Gromov and Whyte in the cubulated case [5, Questions 1.7 and 1.11]. In particular, Theorem A applies to $C^{\prime}(1/6)$ small-cancellation groups [67].
@@
If $G$ is hyperbolic then the result follows from Theorem A. Otherwise, Caprace–Haglund proved that $G$ contains a cubically convex-cocompact subgroup isomorphic to $\mathbb{Z}^{2}$ [13, Corollary 4], and the result follows unless $\mathbb{Z}^{2}$ has finite index. But, in the latter case, $G$ is a surface group by Bieberbach’s theorem. ∎
This corollary answers a question asked by Wise in the 2014 ICM proceedings [68, Question 13.50].
Remark 0.2.
The proof of Corollary B will apply to any class of cubulated groups in which one has a ‘hyperbolisation’ theorem, guaranteeing that a cubulated group $G$ without a $\mathbb{Z}^{2}$ subgroup must be hyperbolic. Such hyperbolisation theorems are also known for cube complexes without cyclic facing triples [51] and for weakly special cube complexes [43].
@@
Historically, Question 0.1 has received the most attention when $G$ is a one-relator group. In this context, the question was attributed to Fine–Rosenberger–Wienke in the Kourovka notebook [35, Question 20.22], but it has also received attention in many other works on one-relator groups, sometimes with additional hypotheses [3, 4, 17, 22]. In this context, it is sometimes referred to as the surface group conjecture. In [65], the author proved the surface group conjecture for Sela’s limit groups, and also for graphs of free groups with cyclic edge groups. [^2] The latter class includes the classes of cyclically pinched and conjugacy pinched one-relator groups considered in [17, 22].
Most recently, the surface group conjecture was addressed by Gardam–Kielak–Logan [23, Conjecture 1.4], who proved it in the case when the one-relator group $G$ has two generators. Combining their work with Theorem A and other recent advances in our understanding of one-relator groups, the surface group conjecture follows.
@@
In fact, the proof of Theorem D produces a finitely generated non-free subgroup unless $G$ has two generators. In the two-generator case, the stronger result is not true: every finitely generated subgroup of infinite index in the $(1,2)$-Baumslag–Solitar group $BS(1,2)$ is either trivial or isomorphic to $\mathbb{Z}$, but $BS(1,2)$ contains a subgroup isomorphic to the infinitely generated non-free group $\mathbb{Q}_{2}$ of dyadic rationals. (Cf. Question 6.4 below.)
@@
Assuming $\pi_{1}(X)$ is one-ended, our next task is to analyse splittings over $\mathbb{Z}$, and to do this we restrict to the case when $\pi_{1}(X)$ is hyperbolic. The approach generalises work of Cashen–Macura [16] and Cashen [15], who showed that a line pattern in a free group does not split over $\mathbb{Z}$ if and only if there is a bound on the diameters of cut pairs in the Whitehead graph. We prove an analogous result, Theorem 3.12, which asserts that $\pi_{1}(X)$ is cyclically indecomposable if and only if there is a uniform bound on the widths of 2-cuts. The point is that cyclic splittings correspond to periodic 2-cuts, from which a pigeonhole argument implies a bound on the width of a 2-cut if $\pi_{1}(X)$ is cyclically indecomposable. Using the projection-complex machinery of Bestvina–Bromberg–Fujiwara [6], and their improved version with Sisto [7], we strengthen this to a uniform bound on the widths of $k$-cuts, independent of $k$.
@@
The paper is structured as follows. Section 1 contains preliminaries on non-positively curved complexes, cube complexes and simplicial complexes. Experts may wish to jump straight to Section 2, where we define Whitehead complexes and use them to analyse free splittings, including proving Theorem E. In Section 3 we extend the analysis to cyclic splittings in the hyperbolic case. Section 4 contains the proofs of the main theorems. Section 5 records a counterexample to a natural generalisation of Theorem A, while some open questions are listed in Section 6.
@@ Acknowledgements
I am grateful to Mark Hagen for useful conversations about this work, most especially for suggesting Lemma 3.20, which greatly simplifies the argument. Thanks also to Macarena Arenas, Jack Button, Chris Cashen, Mark Hagen again, Marco Linton, Lars Louder and Sam Shepherd for comments on a preliminary draft. The second draft benefitted from comments from Martin Bridson, Anthony Genevois and Chris Hruska. Special thanks to Macarena Arenas (again), Misha Kapovich and Genevieve Walsh for help with §5. Finally, thanks to the referees for suggesting many improvements.
@@ 1 Preliminaries
All complexes that we consider will be locally finite.
@@ 1.1 Non-positively curved cube complxes
A cube complex $X$ is a cell complex constructed from a disjoint union of finite-dimensional unit cubes by identifying faces isometrically [10, Example I.7.40(4)]. Further, we will also assume that cubes are embedded, and that the intersection of any two cubes is either empty or a face (i.e. $X$ is cubical in the sense Bridson–Haefliger [10, Definition I.7.32]); this can always be ensured by subdividing.
The link $\mathrm{Lk}_{X}(x)$ of a point $x$ in a metric polyhedral complex $X$ is the space of directions at $x$, i.e. the space of germs of geodesics emanating from $x$ [10, I.7.38], and admits a natural angular metric. When $X$ is a cube complex and $x$ is a vertex, $\mathrm{Lk}_{X}(x)$ has the natural structure of a simplicial complex, where the simplices are the corners of the cubes incident at $x$.
As long as a cube complex $X$ is finite-dimensional, $X$ admits a natural complete geodesic metric in which every $k$-cube is locally isometric to $[0,1]^{k}\subseteq\mathbb{R}^{k}$. We will refer to this metric as the $\ell^{2}$-metric on $X$. Gromov’s link condition asserts that a polyhedral complex is non-positively curved – i.e. the metric is locally CAT(0) – if and only if the angular metric on the link of any point is CAT(1) [27, 4.2.A]. Gromov further proved that, when $X$ is a cube complex, the link condition is satisfied if and only if the link of every vertex is flag, meaning that, for every $k\geq 2$, every $k$-clique in the 1-skeleton of $X$ is the 1-skeleton.
The features of a non-positively curved cube complex $X$ can often be captured by studying its hyperplanes. Let $C$ be a $k$-dimensional cube, identified with $[-1/2,1/2]^{k}\subseteq\mathbb{R}^{k}$. A hypercube of $C$ is an intersection of $C$ with any coordinate hyperplane of $\mathbb{R}^{k}$. The set of hypercubes of cubes in $X$ form a natural directed system under inclusion, and gluing along these inclusions gives a cube complex with a natural map to $X$. The path components of this cube complexes are the hyperplanes of $X$.
@@
Finally, we can make use of the following important lemma [10, Proposition II.4.14].
@@ 1.2 CAT(0) cube complexes
When $X$ is non-positively curved, the natural geodesic metric on the universal cover $\widetilde{X}$ is CAT(0), by the Cartan–Hadamard theorem [27, §4]. This gives us access to the standard tools of CAT(0) geometry detailed in [10], which are especially useful when studying a closed, convex subspace $Y\subseteq\widetilde{X}$. For instance, we may define the orthogonal projection $\pi_{Y}:\widetilde{X}\to Y$ as in [10, Proposition II.2.4].
One important consequence of the uniqueness of geodesics in CAT(0) spaces is that any local isometry between CAT(0) spaces is a global isometric embedding [10, Proposition II.4.14]. Thus, the map from a hyperplane $H\to X$ lifts to a convex embedding of the universal cover $\widetilde{H}\hookrightarrow\widetilde{X}$, and so the hyperplanes of $\widetilde{X}$ are the universal covers of the hyperplanes of $X$.
@@
Fortunately, we may always assume that $X$ is essential by invoking a result of Caprace–Sageev [14, Proposition 3.5].
@@
Note that the $\ell^{1}$-metric on $X^{(1)}$ is also geodesic, and that the $\ell^{1}$ distance between two vertices is exactly the number of hyperplanes that separate them. The next lemma is useful for relating the $\ell^{1}$ and $\ell^{2}$ distances.
@@
Let $Y_{1},Y_{2}$ be convex subcomplexes of a CAT(0) cube complex $\widetilde{X}$. Suppose that a geodesic $\gamma$ minimises the $\ell^{2}$-distance from $Y_{1}$ to $Y_{2}$. A hyperplane $H$ separates $Y_{1}$ from $Y_{2}$ if and only if $\gamma$ crosses $H$.
@@
The necessity is obvious, so it remains to prove that, if $\gamma$ crosses $H$, then $H$ separates $Y_{1}$ from $Y_{2}$. Suppose that $\gamma(0)\in Y_{1}$ and $\gamma$ crosses $H$ at $\gamma(t)$. Because $\gamma$ minimises the distance, it makes an angle of at least $\pi/2$ with each geodesic in $Y_{1}$ through $\gamma(0)$. Hence, if $H$ intersects $Y_{1}$ then $\gamma(0)$, $\gamma(t)$ and $\pi_{H}(\gamma(0))$ form a triangle with two right angles, contradicting the hypothesis that $\widetilde{X}$ is CAT(0). Therefore, $H$ is disjoint from $Y_{1}$, and likewise from $Y_{2}$. Since $\gamma$ crosses $H$ once, it follows that $H$ separates $Y_{1}$ from $Y_{2}$, as required. ∎
We can now state Haglund’s theorem [31, Theorem H]
@@
One useful consequence is to characterise geometrically well-behaved subgroups. A subgroup $H$ of the fundamental group of a non-positively curved cube complex $X$ is called cubically convex-cocompact if $H$ acts cocompactly on a subcomplex $Y\subseteq\widetilde{X}$ that is convex in the $\ell^{2}$-metric.
@@
Let $X$ be a compact, non-positively curved cube complex. A subgroup $H$ of $\pi_{1}(X)$ is cubically convex-cocompact if and only if the $H$-orbit of some vertex is quasiconvex in the $\ell^{1}$-metric.
@@ 1.3 Simplicial complexes
$B$$\mathrm{St_{A}}(B)$$B$$\mathrm{Lk_{A}}(B)$
Figure 1: Open stars and links in simplicial complexes.
@@
Let $A$ be a simplicial complex and $B\subseteq A$ a full subcomplex. The open star of $B$, $\mathrm{St}_{A}(B)$, is the union of the set of simplices of $A$ that contain a simplex of $B$ as a face. The link of $B$, $\mathrm{Lk}_{A}(B)$, is the set of faces of simplices of $\mathrm{St}_{A}(B)$ that are disjoint from $B$. See Figure 1.
@@
In the specific case where $B$ is a set of $k$ pairwise non-adjacent vertices, we say that $B$ is a cut set for $A$. If #$\#B=1$ we say $B$ is a cut vertex, and if #$\#B=2$ we say $B$ is a cut pair. If $A$ has no cut set of cardinality at most $k$, then $A$ is said to be $k$-vertex-connected.
@@ 2 Free splittings
In 1936, J.H.C. Whitehead explained how to determine whether an element $w$ of a free group $F$ is primitive. For a fixed basis $B$ of $F$, his algorithm uses a graph $\mathrm{Wh}_{B}(w)$ which is now known as the Whitehead graph [63]. His algorithm can equally be interpreted as searching for free splittings $F\cong\langle w\rangle*F^{\prime}$. Many elaborations and generalisations of Whitehead’s argument have been given over the years by, among others, Lyon [41], Stallings [59] and Stong [60]. Diao–Feighn, building on work of Gersten [26], gave an algorithm to compute the Grushko decomposition of a graph of free groups [18]. [^3] See the recent preprint of Dicks for a comprehensive history of this topic [19, §2].
The power of this point of view is that global information about a group $G$ – whether or not $G$ is free, or splits freely, or most generally the structure of the Grushko decomposition of $G$ – is encoded in the local information of a Whitehead graph, which can be read off a presentation. The following theorem, which is the result of combining Whitehead’s algorithm with a theorem of Shenitzer [54, Theorem 5], exemplifies this idea in the case of one-relator groups.
@@ 2.1 Bounding hyperplanes and Whitehead complexes
In this section, we introduce the Whitehead complexes that will be our main technical tool. While the use we put them to here appears to be new, they bear some similarities with crossing graphs and crossing complexes for CAT(0) cube complexes, which have been studied extensively in a wide variety of contexts – see, for instance, [24, 28, 45, 49], among many other examples.
@@
The Helly property for hyperplanes of CAT(0) cube complexes, proved by Sageev, asserts that if a set of hyperplanes intersect pairwise then their intersection is non-empty [50, Theorem 4.14]. Hence, Whitehead complexes are flag.
@@
Whitehead complexes are useful because they provide a topological model for the complement of $Y$: we shall see that $\mathrm{Wh}_{X}(Y)$ is homotopy equivalent to the complement $\widetilde{X}\smallsetminus Y$. In order to describe this homotopy equivalence, it is useful to define the link of $Y$, and to extend our discussion of bounding hyperplanes to codimension-$k$ hyperplanes.
@@
The link $\mathrm{Lk}_{\widetilde{X}}(Y)$ is homeomorphic to the realisation of a complex of spaces in the sense of Hatcher, with underlying complex the barycentric subdivision of $\mathrm{Wh}_{\widetilde{X}}(Y)$. Each vertex corresponds to a codimension-$k$ hyperplane $K$, the associated space is the hyperplane component $K_{Y}$, and the attaching maps are given by the natural inclusions. Since each $K_{Y}$ is contractible, it follows that $\mathrm{Lk}_{\widetilde{X}}(Y)\simeq\mathrm{Wh}_{\widetilde{X}}(Y)$ by [33, Proposition 4G.1]. ∎
@@
The lemma can also be deduced from the nerve theorem, by noting that $\mathrm{Wh}_{\widetilde{X}}(Y)$ is the nerve of a natural open cover of $\widetilde{X}\smallsetminus Y$.